OpenStudy (anonymous):

solve using elimination: 6x-15y=39 ; 4x-10y=3

OpenStudy (lncognlto):

6x-15y=39 ...(1) 4x-10y=3 ...(2) (1) x 2: 12x-30y=78 +(2)x(-3): -12x+30y=-9 0=69 It doesn't seem that this equation is solvable?

OpenStudy (anonymous):

no solution, inconsistent?

OpenStudy (lncognlto):

Yep, it's inconsistent.